The U. S- based MNC Conte has receivables of JPY 5 million in three months. The spot JPYUSD is 0.0085. The firm’s economist forecasts that the JPYUSD could end the period with a value of either 0.00825 (probability of 45 percent) or 0.00875 (55 percent). The firm is concerned about its currency risk. It has also assessed some hedging alternatives. Three- month JPY forward contracts are traded at USD 0.00855. The three- month interest rates (annual compounding) in the United States and Japan are 3 percent and 0.5 percent, respectively. Put options with three- month expiration and a strike price of USD 0.00850 are available for a premium of USD 0.00032. What is the expected cash flow (and the corresponding standard deviation) if company uses money market hedge?
Answer #1

Expected JPYUSD at the end of the period = (0.45×0.00825)+(0.55×0.00875) = 0.008525

Steps for Money Market Hedge:

JPY 5 Million are receivable. Therefore, liability in JPY should be created.

Now,

(1) Borrow [5,000,000/(1+0.005/4)] = ¥4,993,758 for 3 months @0.5% p.a.

(2) Conert ¥4,993,758 into $ at spot rate i.e. 0.0085 and receive $42,446.943

(3) Invest $42,446,943 for 3 months @3% p.a.

After 3 months,

(4) Receive ¥5,000,000 that are receivable from exports.

(5) Repay borrowed loan alongwith interest i.e. ¥4,993,758 + 0.125% = ¥5,000,000

(6) Receive proceeds from investment i.e. $42,446.943 + 0.75% = $42,765.295

Therefore, Expected Cash Flow by using Money Market Hedge will be $42,765.295

There will be NO STANDARD DEVIATION in using Money Market Hedge. This cash flow is SURE.

Also, HEDGING is done to ELIMINATE THE VARIATION RISK. So, there will not be any deviation in that.